OpenStudy (nuccioreggie):

Ross needs to buy a countertop for a laundry room. He calculated the area to be 12 square feet. The actual area is 11.8 square feet. What is Ross's percentage error? 1.02% 1.69% 20% 98%

OpenStudy (cwrw238):

%age error = difference in actual and calculated * 100 ----------------------------------- actual area

OpenStudy (nuccioreggie):

So its a @cwrw238

OpenStudy (anonymous):

All we need to do is to find how many percent of 12 is 0.2 (why 0.2? because that's his error which can be counted by 12 - 11.8 = 0.2). 0.2 = x% of 12 0.2 = x/100 * 12 0.2 = 12x / 100 / * 100 (both sides) 20 = 12x / ÷ 12 (both sides) x = 20/12 x = 5/3 x = 1 2/3 x = 1.66 Answer: Ross' percentage error was about 1.66%.

OpenStudy (cwrw238):

no try again (12 - 11.8) * 100 -------------- 11.8

OpenStudy (nuccioreggie):

Oooooooo Ok so its B

OpenStudy (cwrw238):

1.69 B

OpenStudy (nuccioreggie):

Correct?

OpenStudy (cwrw238):

yes

OpenStudy (cwrw238):

Faamous 5 was wrong- the percentage is based on the actual area which was 11.8 not 12